OpenStudy (anonymous):

ssume that the terminal side of an angle of t radians in standard position lies in the given quadrant on the given straight line. Find sin (t), cos (t), tan (t). Quadrant III; line through the origin parallel to: 4x - 9y = -8

OpenStudy (anonymous):

I'm guessing it is on 4x-9y=0

OpenStudy (anonymous):

hmm... idk why but it took me a while to understand this problem...

OpenStudy (anonymous):

do you know the equation of the line through the origin parallel to 4x - 9y = -8?

OpenStudy (anonymous):

Well, I know it has to be 4x-9y=0. It's in Quadrant 3, so I think both (x,y) are going to be negative.

OpenStudy (anonymous):

I just really don't know where to go from there.

OpenStudy (anonymous):

Well, I know I need to find a radius, but I don't know what points to put in.

OpenStudy (anonymous):

ok... so here's a sketch of the line along with angle t....